# EnKF-C user guide version 2.32.0

### v11

@@ 2.4.4 Observation types parameter file
     >./bin/enkf_prep --describe-prm-format obstypes      Observation types parameter file format:        NAME        = <name>    [ DOMAINS     = <domain name> ... ]      ISSURFACE   = {yes | no}    [ STATSONLY   = {yes | no*} ]      VAR         = <model variable name> ...    [ ALIAS       = <variable name used in file names> ]   (VAR*)    [ OFFSET      = <file name> <variable name> ]          (none*)    [ MLD_VARNAME = <model varname> ]                      (none*)    [ MLD_THRESH  = <threshold> ]                          (NaN*)    [ HFUNCTION   = <H function name> ]                    (standard*)    [ ASYNC       = <time interval> [centre*|endpoint [time varname]]] (0*)    [ LOCRAD      = <loc. radius in km> ... ]    [ LOCWEIGHT   = <loc. weight> ... ]                    (# LOCRAD > 1)    [ RFACTOR     = <rfactor> ]                            (1*)    [ NLOBSMAX    = <max. allowed number of local obs.> ]  (inf*)    [ ERROR_STD_MIN = <min. allowed superob error> ]       (0*)    [ SOBSTRIDE   = <stride for superobing> ]              (1*)    [ PERMIT_LOCATION_BASED_THINNING = <yes | no> ]        (yes*)    [ MINVALUE    = <minimal allowed value> ]              (-inf*)    [ MAXVALUE    = <maximal allowed value> ]              (+inf*)    [ XMIN        = <minimal allowed X coordinate> ]       (-inf*)    [ XMAX        = <maximal allowed X coordinate> ]       (+inf*)    [ YMIN        = <minimal allowed Y coordinate> ]       (-inf*)    [ YMAX        = <maximal allowed Y coordinate> ]       (+inf*)    [ ZMIN        = <minimal allowed Z coordinate> ]       (-inf*)    [ ZMAX        = <maximal allowed Z coordinate> ]       (+inf*)    [ WINDOWMIN   = <start of obs window in days from analysis> ] (-inf*)    [ WINDOWMAX   = <end of obs window in days from analysis> ]   (+inf*)      [ <more of the above blocks> ]
@@ 2.4.5 Observation data parameter file
     ./bin/enkf_prep --describe-prm-format obsdata      Observation data parameter file format:        PRODUCT   = <product>      READER    = <reader>    [ PARAMETER <name> = <value> ]      ...      TYPE      = <observation type>      FILE      = <data file wildcard>      ...    [ ERROR_STD = { <value> | <data file> <varname> } [ EQ* | PL | MU | MI | MA ] ]      ...      [ <more of the above blocks> ]      [ EXCLUDE   = { <observation type> | ALL } <lon1> <lon2> <lat1> <lat2> ]      ...     Observation files can be defined using wildcards “*” and “?”. Missing a file is reported in the log and is not considered to be a fatal error.
@@
     # set observation error for Geosat to 7cm  PRODUCT = RADS  TYPE = SLA  READER = scattered  PARAMETER VARNAME = sla  PARAMETER ZVALUE = 0  PARAMETER MINDEPTH = 100  FILE = /short/p93/pxs599/obs/RADS/2006/g?_20060523.nc  ERROR_STD = 0.07    # use default errors for other altimeters  PRODUCT = RADS  TYPE = SLA  READER = scattered  PARAMETER VARNAME = sla  PARAMETER ZVALUE = 0  PARAMETER MINDEPTH = 100  file = /short/p93/pxs599/obs/RADS/2006/[!g]?_20060523.nc
@@ 2.5 File name conventions
[…]
@@ 2.6 PREP
     >./bin/enkf_prep    Usage: enkf_prep <prm file> [<options>]    Options:    --allow-logspace-with-static-ens        confirm that static ensemble is conditioned for using log space    --consider-subgrid-variability        increase error of superobservations according to subgrid variability    --describe-prm-format [main|model|grid|obstypes|obsdata]        describe format of a parameter file and exit    --describe-reader <reader>        decribe reader and exit    --describe-superob <sob #>        print composition of this superobservation and exit    --list-readers        list available readers and exit    --no-superobing    --no-superobing-across-batches    --no-thinning    --superob-across-instruments    --write-orig-obs        write original obs within model extent to observations-orig.nc    --write-all-orig-obs        write all original obs to observations-orig.nc    --version        print version and exit
@@ 2.6.1 Observation types, products, instruments, batches, readers
     PRODUCT = RADS  TYPE = SLA  READER = scattered  PARAMETER VARNAME = sla  PARAMETER BATCHNAME = pass  PARAMETER ZVALUE = 0  PARAMETER MINDEPTH = 100  FILE = obs/RADS/2007/??_200712{19,20,21,22,23}.nc    PRODUCT = ESACCI  TYPE = SST  READER = scattered  PARAMETER VARNAME = sst  PARAMETER ZVALUE = 0  PARAMETER VARSHIFT = -273.15  FILE = obs/ESACCI/2007/200712{19,20,21,22,23}-*.nc
@@
>bin/enkf_prep --list-readers  generic readers:    scattered    gridded_xy    gridded_xyz    gridded_xyh    z  custom readers:    navo    windsat    mmt    amsr2    amsre    cmems
Users are encouraged to use generic readers. When this is not possible, one may either modify (or put a request to modify) a generic reader, or develop a custom reader.
Each reader can be specified in the observation data parameter file with an arbitrary number of parameters. For example, the following section changes the default minimal depth for using altimetry observations to 150 m:
(...)PRODUCT == RADSTYPE = SLAREADER = scatteredPARAMETER VARNAME = slaPARAMETER ZVALUE = 0PARAMETER MINDEPTH = 150(...)
@@
>./bin/enkf_prep --describe-reader scattered  Generic reader "scattered" reads 2D or 3D scattered point data.  There are a number of parameters that must (marked below with "++"), can  ("+"), or may ("-") be specified in the corresponding section of the  observation data parameter file. The names in brackets represent the default  names checked in the abscence of the entry for the parameter. Each parameter  needs to be entered as follows:    PARAMETER <name> = <value> ...  Parameters common to generic readers:    - VARNAME (++)    - TIMENAME ("*[tT][iI][mM][eE]*") (+)    - or TIMENAMES (when time = base_time + offset) (+)    - LONNAME ("lon" | "longitude") (+)    - LATNAME ("lat" | "latitude") (+)    - ZNAME ("z") | ZVALUE (+)        "ZNAME" is needed for 3D data, "ZVALUE" for 2D data (can be NaN)    - STDNAME ("std") (-)        dispersion of the collated data    - ESTDNAME ("error_std") (-)        error STD; if absent then needs to be specified in the corresponding        section of the observation data parameter file    - BATCHNAME ("batch") (-)        name of the variable used for batch ID (e.g. "pass" for SLA)    - INSTRUMENT (-)        instrument string that will be used for calculating instrument stats        (overrides the global attribute "instrument" in the data file)    - QCFLAGNAME (-)        name of the QC flag variable, possible values 0 <= qcflag <= 31    - QCFLAGVALS (-)        the list of allowed values of QC flag variable        Note: it is possible to have multiple entries of QCFLAGNAME and        QCFLAGVALS combination, e.g.:          PARAMETER QCFLAGNAME = TEMP_quality_control          PARAMETER QCFLAGVALS = 1          PARAMETER QCFLAGNAME = DEPTH_quality_control          PARAMETER QCFLAGVALS = 1          PARAMETER QCFLAGNAME = LONGITUDE_quality_control          PARAMETER QCFLAGVALS = 1,8          PARAMETER QCFLAGNAME = LATITUDE_quality_control          PARAMETER QCFLAGVALS = 1,8        An observation is considered valid if each of the specified flags takes        a permitted value.  Parameters specific to the reader:    - ADDVAR (-)        name of the variable to be added to the main variable (can be repeated)    - SUBVAR (-)        name of the variable to be subtracted from the main variable (can be        repeated)  Parameters common to all readers:    - VARSCALE (-)        scale factor (applied before VARSHIFT)    - VARSHIFT (-)        data offset to be added (e.g. -273.15 to convert from K to C)    - FOOTRPINT (-)        footprint of observations in km    - MINDEPTH (-)        minimal allowed model depth    - MAXDEPTH (-)        maximal allowed model depth    - STRIDE (-)        stride interval.    - YMIN (-)        minimal allowed latitude    - YMAX (-)        maximal allowed latitude    - ZMIN (-)        minimal allowed depth    - ZMAX (-)        maximal allowed depth
@@ 2.6.2 Superobing
The observations are merged by averaging their values, coordinates and times with weights inversely proportional to the observation error variance. The observation error variance of a superobservation is set to the inverse of the sum of inverse observation error variances of the merged observations. The product and instrument fields of the superobservation are set either to those of the merged observations or to -1, depending on whether the merged observations have the same values for these fields or not.
@@
Note that during superobing EnKF-C by default thins observations with identical positions, assuming that those must been obtained from high-frequency instruments (e.g. moorings). This thinning can be switched off by the command line parameter --no-thinning, or for observations of a particular data type only by adding flag PERMIT_LOCATION_BASED_THINNING = no to the corresponding section in the observation types parameter file.
@@ 2.6.3 Asynchronous DA / FGAT
[…]
       calculating ensemble observations:    2014-03-22 06:28:28      ensemble size = 96      distributing iterations:        all processes get 6 iterations        process 0: 0 - 5      SST |aaaaaa|aaaaaa|aaaaaa|aaaaaa|aaaaaa      SLA |aaaaaa|aaaaaa|aaaaaa|aaaaaa|aaaaaa      TEM ......      SAL ......     The entries “a” mean that the observations are assimilated asynchronously and the files and the corresponding (by name) files have been found. These entries would be replaced by “s” if the observations were assimilated synchronously because of lacking the corresponding files. The vertical lines indicate the time slots for asynchronous DA; in the above example the DAW has 5 time slots. The entries “.” indicate calculating ensemble observations for synchronous observations. Note that only the master process is writing to the log here, which explains why there is only output from 6 members in the log above.
@@
Note that if the model dump associated with a particular time bin is not found, then the corresponding restart file is used; i.e. the observations from this time bin are assimilated sychronously. This behaviour can be disallowed by running CALC with option --strict-time-matching. Using this option requires specifying the name of the time variable in model dumps with the entry ASYNC of the observation types parameter file. It is strongly encouraged to use this option when possible. The strict time matching permits skipping the model dump associated with a time bin only if there is a single time bin for the observation type, and its centre matches the time of the restart dump.
@@ 2.7.1 Observation functions
     NAME = SLA  ...  HFUNCTION = standard  ...     The available functions for each observation type are specified by the variable allhentries in calc/allhs.c. The “standard” functions do normally perform 2D or 3D linear interpolation from the corner model grid nodes for the cell containing the observation.
@@ 2.7.2 Interpolation of ensemble transforms
Local ensemble transforms $\mathbf{X}_{5}$ (EnKF) or local ensemble weights $\mathbf{w}$ (EnOI) represent smooth fields with the characteristic spatial variability scale of the localisation radius. This smoothness allows one to reduce the computational load in CALC by calculating local transforms or weights on a subgrid with a specified stride only, and using linearly interpolated transforms or weights in the intermediate grid cells (Yang et al., 2009). The value of the stride is defined by the STRIDE entry in the main parameter file and can be overwritten for a particular grid in the grid parameter file.

### v12

@@ 2.4.4 Observation types parameter file
      >./bin/enkf_prep --describe-prm-format obstypes         Observation types parameter file format:         NAME = <name>     [ DOMAINS = <domain name> ... ]     ISSURFACE = {yes | no}     [ STATSONLY = {yes | no*} ]     VAR = <model variable name> ...     [ ALIAS = <variable name used in file names> ] (VAR*)     [ OFFSET = <file name> <variable name> ] (none*)     [ MLD_VARNAME = <model varname> ] (none*)     [ MLD_THRESH = <threshold> ] (NaN*)     [ HFUNCTION = <H function name> ] (standard*)     [ ASYNC = <time interval> [{centre* | endpoint} [time varname]]] (0*)     [ LOCRAD = <loc. radius in km> ... ]     [ LOCWEIGHT = <loc. weight> ... ] (# LOCRAD > 1)     [ RFACTOR = <rfactor> ] (1*)     [ ERROR_DOUBLING_TIME = <time in days> ] (inf*)     [ NLOBSMAX = <max. allowed number of local obs.> ] (inf*)     [ ERROR_STD_MIN = <min. allowed superob error> ] (0*)     [ SOBSTRIDE = <stride for superobing> ] (1*)     [ PERMIT_LOCATION_BASED_THINNING = {yes* | no} ]     [ MINVALUE = <minimal allowed value> ] (-inf*)     [ MAXVALUE = <maximal allowed value> ] (+inf*)     [ XMIN = <minimal allowed X coordinate> ] (-inf*)     [ XMAX = <maximal allowed X coordinate> ] (+inf*)     [ YMIN = <minimal allowed Y coordinate> ] (-inf*)     [ YMAX = <maximal allowed Y coordinate> ] (+inf*)     [ ZMIN = <minimal allowed Z coordinate> ] (-inf*)     [ ZMAX = <maximal allowed Z coordinate> ] (+inf*)     [ WINDOWMIN = <start of obs window in days from analysis> ] (-inf*)     [ WINDOWMAX = <end of obs window in days from analysis> ] (+inf*)         [ <more of the above blocks> ]
@@ 2.4.5 Observation data parameter file
      ./bin/enkf_prep --describe-prm-format obsdata         Observation data parameter file format:         PRODUCT = <product>     TYPE = <observation type>     READER = <reader>     FILE = <data file wildcard>     ...     [ ERROR_STD = { <value> | <data file> <varname> } [ EQ* | PL | MU | MI | MA ] ]     ...     [ MANDATORY = {yes | no*} ]     [ PARAMETER <name> = <value> ]     ...         [ <more of the above blocks> ]         [ EXCLUDE = { <observation type> | ALL } <lon1> <lon2> <lat1> <lat2> ]     ...      Observation files can be defined using wildcards “*” and “?”. Missing a file is reported in the log and is not considered to be a fatal error.
@@
      # set observation error for Geosat to 7cm    PRODUCT = RADS    TYPE = SLA    READER = scattered    PARAMETER VARNAME = sla    PARAMETER ZVALUE = NaN    PARAMETER MINDEPTH = 100    FILE = /short/p93/pxs599/obs/RADS/2006/g?_20060523.nc    ERROR_STD = 0.07        # use default errors for other altimeters    PRODUCT = RADS    TYPE = SLA    READER = scattered    PARAMETER VARNAME = sla    PARAMETER ZVALUE = NaN    PARAMETER MINDEPTH = 100    file = /short/p93/pxs599/obs/RADS/2006/[!g]?_20060523.nc
@@ 2.5 File name conventions
EnKF-C assumes that the ensemble and background file names have some predefined formats. The file name for ensemble member memberid containing model variable varname is assumed to be sprintf("mem%03d_%s.nc", memberid, varname). This file typically represents the model restart (EnKF) or ensemble anomaly (EnOI).
For the EnOI the background (forecast restart) file for variable varname is assumed to be sprintf("bg_%s.nc", varname).
The analysis restart file names are created by concatenating the above forecast restart names and a suffix, either“.analysis” or “.increment”, depending on the output of UPDATE.
For asynchronous DA the model output file names for the time slot t are assumed to be
sprintf("mem%03d_%s_%d.nc", memberid, varname, t) (EnKF) and sprintf("bg_%s_%d.nc", varname, t) (EnOI). It is also possible to have model outputs concatenated into a single multi-record file, in which case it is assumed to be either sprintf("mem%03d_%s_#.nc", memberid, varname) (EnKF) or sprintf("bg_%s_#.nc", varname). In this case EnKF-C will identify the record to be used by its time. (See sec. 2.6.3.)
@@ 2.6 PREP
      >./bin/enkf_prep     Usage: enkf_prep <prm file> [<options>]     Options:     --consider-subgrid-variability     increase error of superobservations according to subgrid variability     --describe-prm-format [main|model|grid|obstypes|obsdata]     describe format of a parameter file and exit     --describe-reader <reader>     decribe reader and exit     --describe-superob <sob #>     print composition of this superobservation and exit     --list-readers     list available readers and exit     --no-superobing     --no-superobing-across-batches     --no-thinning     --superob-across-instruments     --write-orig-obs     write original obs within model extent to observations-orig.nc     --write-all-orig-obs     write all original obs to observations-orig.nc     --version     print version and exit
@@ 2.6.1 Observation types, products, instruments, batches, readers
      PRODUCT = RADS    TYPE = SLA    READER = scattered    PARAMETER VARNAME = sla    PARAMETER BATCHNAME = pass    PARAMETER ZVALUE = NaN    PARAMETER MINDEPTH = 100    FILE = obs/RADS/2007/??_200712{19,20,21,22,23}.nc        PRODUCT = ESACCI    TYPE = SST    READER = scattered    PARAMETER VARNAME = sst    PARAMETER ZVALUE = 0    PARAMETER VARSHIFT = -273.15    FILE = obs/ESACCI/2007/200712{19,20,21,22,23}-*.nc
@@
>bin/enkf_prep --list-readers generic readers: scattered gridded_xy gridded_xyz gridded_xyh profile custom readers: navo windsat mmt amsr2 amsre cmems
Users are encouraged to use generic rather than custom readers. When this is not possible, one may either modify (or put a request to modify) a generic reader, or develop a custom reader.
Each reader can be specified in the observation data parameter file with an arbitrary number of parameters. For example, the following section sets the default minimal depth for SLA observations to 150 m:
(...)PRODUCT == RADSTYPE = SLAREADER = scatteredPARAMETER VARNAME = slaPARAMETER ZVALUE = NaNPARAMETER MINDEPTH = 150(...)
@@
LOCATION_BASED_THINNING_TYPE – type of location based thinning (XYZ* | XY | CELL | NILL).
Specified parameters apply to observations defined in the corresponding section of observation data parameter file.
@@
>./bin/enkf_prep --describe-reader scattered Generic reader "scattered" reads 2D or 3D scattered point data. There are a number of parameters that must (marked below with "++"), can ("+"), or may ("-") be specified in the corresponding section of the observation data parameter file. The names in brackets represent the default names checked in the abscence of the entry for the parameter. Each parameter needs to be entered as follows: PARAMETER <name> = <value> ... Parameters common to generic readers: - VARNAME (++) - TIMENAME ("*[tT][iI][mM][eE]*") (+) - or TIMENAMES (when time = base_time + offset) (+) - LONNAME ("lon" | "longitude") (+) - LATNAME ("lat" | "latitude") (+) - ZNAME ("z") | ZVALUE (+) "ZNAME" is needed for 3D data, "ZVALUE" for 2D data (can be NaN) - STDNAME ("std") (-) dispersion of the collated data - ESTDNAME ("error_std") (-) error STD; if absent then needs to be specified in the corresponding section of the observation data parameter file - BATCHNAME ("batch") (-) name of the variable used for batch ID (e.g. "pass" for SLA) - INSTRUMENT (-) instrument string that will be used for calculating instrument stats (overrides the global attribute "instrument" in the data file) - QCFLAGNAME (-) name of the QC flag variable, possible values 0 <= qcflag <= 31 - QCFLAGVALS (-) the list of allowed values of QC flag variable Note: it is possible to have multiple entries of QCFLAGNAME and QCFLAGVALS combination, e.g.: PARAMETER QCFLAGNAME = TEMP_quality_control PARAMETER QCFLAGVALS = 1 PARAMETER QCFLAGNAME = DEPTH_quality_control PARAMETER QCFLAGVALS = 1 PARAMETER QCFLAGNAME = LONGITUDE_quality_control PARAMETER QCFLAGVALS = 1,8 PARAMETER QCFLAGNAME = LATITUDE_quality_control PARAMETER QCFLAGVALS = 1,8 An observation is considered valid if each of the specified flags takes a permitted value. Parameters specific to the reader: - ADDVAR (-) name of the variable to be added to the main variable (can be repeated) - SUBVAR (-) name of the variable to be subtracted from the main variable (can be repeated) Parameters common to all readers: - VARSCALE (-) scale factor (applied before VARSHIFT) - VARSHIFT (-) data offset to be added (e.g. -273.15 to convert from K to C) - FOOTRPINT (-) footprint of observations in km - MINDEPTH (-) minimal allowed model depth - MAXDEPTH (-) maximal allowed model depth - STRIDE (-) stride interval. - YMIN (-) minimal allowed latitude - YMAX (-) maximal allowed latitude - ZMIN (-) minimal allowed depth - ZMAX (-) maximal allowed depth - LOCATION_BASED_THINNING_TYPE XYZ* | XY | CELL | NIL
@@ 2.6.2 Superobing
The observations are merged by averaging their values, coordinates and times with weights inversely proportional to the observation error variance. The observation error variance of a superobservation is set to the inverse of the sum of inverse observation error variances of the merged observations. The product and instrument fields of the superobservation are set either to those of the merged observations or to -1, depending on whether the merged observations have identical values for these fields or not.
@@
Note that during superobing EnKF-C thins observations with identical positions, assuming that those must been obtained from high-frequency instruments (e.g. moorings). The thinning involves replacing the corresponding batch of observations by batch average values. The type of thinning can be set by parameter LOCATION_BASED_THINNING_TYPE in observation data parameter file and can be set to “XYZ” (thinning observations with coinciding X,Y,Z coordinates), “XY” (coinciding X,Y coordinates only), “CELL” (all observations within cell), or “NILL” (no thinning). The location based thinning can be switched off by the command line parameter --no-thinning, or for observations of a particular data type only by adding flag PERMIT_LOCATION_BASED_THINNING = no in the corresponding section in the observation types parameter file.
@@ 2.6.3 Asynchronous DA / FGAT
The model dumps for each asynchronous interval are read from files with names mem<xxx>_<variable name>_<time shift>.nc in the ensemble directory (for the EnKF) or bg_<variable name>_<time shift>.nc in the background directory (for the EnOI). Here “time shift” is the interval ID (with the interval 0 being centred/starting at the observation time). If the corresponding members (or the background files, in the case of EnOI) are found, the observations are assimilated asynchronously; if they are not found, then the observations are assimilated synchronously. This can be tracked from the CALC log file, e.g.:
       calculating ensemble observations:     2014-03-22 06:28:28     ensemble size = 96     distributing iterations:     all processes get 6 iterations     process 0: 0 - 5     SST |aaaaaa|aaaaaa|aaaaaa|aaaaaa|aaaaaa     SLA |aaaaaa|aaaaaa|aaaaaa|aaaaaa|aaaaaa     TEM ......     SAL ......      The entries “a” mean that the observations are assimilated asynchronously and the corresponding files have been found. These entries would be replaced by “s” if the observations were assimilated synchronously because of lacking the corresponding files. The vertical lines indicate the time slots for asynchronous DA; in the above example the DAW has 5 time slots. The entries “.” indicate calculating ensemble observations for synchronous observations. Note that only the master process is writing to the log here, which explains why there is only output from 6 members in the log above.
@@
Note that if the model dump associated with a particular time bin is not found, then the corresponding restart file is used; i.e. the observations from this time bin are assimilated synchronously. This behaviour can be disallowed by running CALC with option --strict-time-matching. Using this option requires specifying the name of the time variable in model dumps with the entry ASYNC of the observation types parameter file. It is strongly encouraged to use this option when possible. The strict time matching permits skipping the model dump associated with a time bin only if there is a single time bin for the observation type, and its centre matches the time of the restart dump.
From version 2.36.0 it is possible to consolidate model dumps used for calculating forecast observations for an observation type for each time interval in the observation window in a single file with multiple time records. Such multi-record files should have name mem<xxx>_<variable name>_#.nc. For each time binning interval the code finds and uses the record with time matching the interval time.
@@ 2.7.1 Observation functions
      NAME = SLA    ...    HFUNCTION = standard    ...      The available functions for each observation type are specified by the variable allhentries in calc/allhs.c:
typedef struct { int issurface; char* H_tag; H_fn H;} H_entry;H_entry allhentries[] = { {1, "standard", H_surf_standard}, {1, "biased", H_surf_biased}, {0, "standard", H_subsurf_standard}, {0, "wsurfbias", H_subsurf_wsurfbias}, {0, "lowmem", H_subsurf_lowmem}, /* * (the corresponding observation type can be entered either as * "issurface = 1" or "issurface = 0") */ {1, "vertsum", H_vertsum}, {0, "vertsum", H_vertsum}, {1, "vertwavg", H_vertwavg}, {0, "vertwavg", H_vertwavg}};
H-functions “standard” are used by default. They perform 2D or 3D linear interpolation on the model grid associated with the observation type.
Function “lowmem” can be used instead of “normal” for observation types associated with 3D model variables in situations with very large models. Unlike “normal” it does not read the whole 3D dump for a variable but proceeds by reading only two layers at a time.
Function “biased” targets 2D variables corrected for bias. Function “wsurfbias” was developed to offset the values of temperature in the mixed layer by the value of SST bias.
Function “vertsum” is used to assimilate observations corresponding to the sum of a model variable over all layers (such as the total sea-ice concentration). It can be used only for variables on sigma or hybrid grids. Function “vertwavg” matches observations with a weighted sum of a model variable in all model layers.
@@ 2.7.2 Interpolation of ensemble transforms
Local ensemble transforms $\mathbf{X}_{5}$ (EnKF) or ensemble weights $\mathbf{w}$ (EnOI) (sec. 1.3.1) represent smooth fields with the characteristic spatial variability scale of the localisation radius. This makes it possible to reduce the computational load in CALC by calculating local transforms or weights on a subgrid with specified stride and then linearly interpolating the transforms or weights in the intermediate grid cells (Yang et al., 2009). The value of the stride is defined by the entry STRIDE in the main parameter file and can be overwritten for a particular grid in the grid parameter file.